### v1

# Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech

Abstract
Integrating human feedback to align text-to-speech (TTS) system outputs with human preferences has proven to be an effective approach for enhancing the robustness of language model-based TTS systems. Current approaches primarily focus on using preference data annotated at the utterance level. However, frequent issues that affect the listening experience often only arise in specific segments of audio samples, while other segments are well-generated. In this study, we propose a fine-grained preference optimization approach (FPO) to enhance the robustness of TTS systems. FPO focuses on addressing localized issues in generated samples rather than uniformly optimizing the entire utterance. Specifically, we first analyze the types of issues in generated samples, categorize them into two groups, and propose a selective training loss strategy to optimize preferences based on fine-grained labels for each issue type. Experimental results show that FPO enhances the robustness of zero-shot TTS systems by effectively addressing local issues, significantly reducing the bad case ratio, and improving intelligibility. Furthermore, FPO exhibits superior data efficiency compared with baseline systems, achieving similar performance with fewer training samples.
Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech
Jixun Yao1, Yuguang Yang2, Yu Pan2, Yuan Feng2, Ziqian Ning1, Jianhao Ye2, Hongbin Zhou2, Lei Xie1 1Northwestern Polytechnical University 2Everest Team, Ximalaya
1 Introduction
Large language models (LLMs) demonstrate remarkable in-context learning abilities in natural language processing (NLP) tasks by scaling both model parameters and dataset size (Brown et al., 2020; Touvron et al., 2023; Grattafiori et al., 2024). Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017) is crucial for calibrating these models, aligning them with human preferences, and minimizing the risk of generating harmful or toxic content. This technique is essential for the success of systems like ChatGPT (Achiam et al., 2023) and has recently been extended to tasks in other modalities.
Speech, a key modality of generative artificial intelligence, is experiencing rapid advancements inspired by developments in text-based LLMs. Recent text-to-speech (TTS) studies (Wang et al., 2023; Chen et al., 2024b; Du et al., 2024) tokenize speech signals into discrete units and employ language model to generate speech tokens in a next-token prediction manner based on phoneme inputs and prompt speech recordings. By leveraging in-context learning capabilities, these TTS systems can produce high-quality, personalized speech in a zero-shot manner. Similar to the use of RLHF in calibrating text LLMs, there is growing interest in integrating RLHF into TTS systems (Zhang et al., 2024; Chen et al., 2024a; Hu et al., 2024; Tian et al., 2024). These approaches effectively enhance the robustness and zero-shot capabilities of TTS systems by incorporating human evaluation.
Despite significant advancements in integrating RLHF into TTS systems, we believe two primary challenges remain. The first challenge is efficiently calibrating pre-trained TTS systems with limited preference data. RLHF typically requires preference data evaluated by human annotators, which is then used for training to align TTS systems with human preferences. However, obtaining large-scale human-annotated speech preference data is prohibitively expensive. Compared to annotating text data, annotating speech data is more difficult due to its inclusion of both linguistic content and paralinguistic information.
The second challenge is achieving fine-grained alignment of specific segments in generated samples with human feedback. In recent years, issues such as poor speaker similarity and degraded audio quality, referred to as systemic errors, have been largely mitigated in TTS systems through advanced technology. However, more frequent issues that impact the listening experience often only occur in specific segments of the audio samples, which we call segmental errors. These include issues like mispronunciation of certain words, abnormal pauses, or repetition, etc. However, current speech RLHF approaches typically annotate preference data at the utterance level, which fails to account for fine-grained issues. Additionally, utterance-level preference optimization calculates the loss for the entire sequence, even though many tokens are already well-trained, leading to inefficiencies (Lin et al., 2024b). We believe that optimizing problematic segments in a fine-grained manner would be both more effective and more data-efficient.
To address the above challenges, we propose a novel fine-grained preference optimization (FPO) framework—an RLHF-based approach aimed at addressing segmental errors in generated samples and improving the robustness of zero-shot TTS systems. FPO employs a fine-grained sampling-annotating pipeline, where representative audio samples are first generated multiple times by TTS systems. Preference data are then annotated from human listening perspectives, labeling each sample as preferred or dispreferred at the utterance level. For each paired preference data, we annotate the problematic segment of the local issue to obtain fine-grained token-level preference labels, enabling more precise optimization than traditional approaches. Furthermore, we analyze the types of issues across different problem segments from an evaluation perspective and propose a selective training loss strategy to optimize each type of issue. During training, FPO processes the entire sequence through the TTS models but computes loss exclusively for the problematic segments, which can effectively resolve fine-grained issues while enhancing training efficiency. Experimental results demonstrate that FPO significantly enhances the robustness of zero-shot TTS systems and exhibits superior data efficiency compared with baseline systems, improving the quality of generated speech across both subjective and objective metrics. The main contributions can be summarized as follows:
  • We propose FPO, an RLHF-based approach designed to align problematic segments in generated samples, thereby enhancing the robustness of TTS systems. The key novelty of FPO lies in its preference optimization strategy, which focuses on optimizing specific problematic segments while avoiding unnecessary computation on well-learned tokens, resulting in a more data-efficient optimization process.
  • By analyzing issues in generated samples, FPO fine-grained optimizes TTS systems on problematic segments using a selective training strategy. This approach introduces a new perspective for addressing fine-grained fixes in TTS preference optimization, focusing on resolving bad cases efficiently and effectively.
  • Extensive experiments demonstrate the efficiency and effectiveness of FPO across both subjective and objective metrics. Notably, FPO significantly reduces the occurrence of bad cases compared to powerful baseline systems, effectively addressing the robustness challenges of zero-shot TTS systems. Audio samples can be found at https://yaoxunji.github.io/fpo.
2 Related Work
2.1 Neural codec based Zero-shot TTS
Inspired by the success of LLMs in NLP, formatting the TTS task as a next-token prediction problem has gained significant popularity in recent years (Borsos et al., 2022; Zhang et al., 2023). VALL-E (Wang et al., 2023), as a pioneering work, introduced the first neural codec based TTS framework, leveraging large-scale speech data to develop in-context learning capabilities. It can synthesize high-quality personalized speech in a zero-shot manner using only a 3-second enrolled recording as an acoustic prompt. Building on this foundation, the typical neural codec-based TTS systems (Kharitonov et al., 2023; Xin et al., 2024; Peng et al., 2024) involve tokenizing speech signals into discrete tokens using a neural codec (Défossez et al., 2022; Zeghidour et al., 2021). A decoder-only language model is then trained to predict these discrete tokens alongside phoneme sequences. Finally, the predicted tokens are reconstructed into speech signals using the codec decoder. This approach has demonstrated promising scalability with large datasets and model sizes, enabling high-quality synthesized speech and powerful zero-shot capabilities for unseen speakers (Łajszczak et al., 2024; Du et al., 2024).
Figure 1: Typical examples of segmental errors can be categorized into two types: temporal modeling errors and semantic-phonetic alignment errors. The problematic segments in the generated samples are highlighted using red boxes or symbols.
2.2 Reinforcement Learning from Human Feedback
Human feedback has been widely used in language models for NLP tasks (Ouyang et al., 2022), with RLHF technology playing a crucial role in aligning models with human preferences. Conventionally, preference optimization is performed by maximizing rewards computed from a reward model to align generated content. Recent approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), leverage closed-form loss functions that directly operate on preference data. This method eliminates the need for explicit reward model training while achieving alignment performance comparable to traditional RLHF.
RLHF in TTS remains in its infancy compared to its advancements in NLP. SpeechAlign first introduced RLHF in TTS (Zhang et al., 2024), exploring multiple preference alignment methods using paired preference data. UNO (Chen et al., 2024a) and RIO (Hu et al., 2024) extended this by optimizing unpaired preference data: UNO accounts for annotation uncertainty in subjective evaluations, while RIO introduces a reverse preference data selection method based on Bayesian principles. Recent studies (Tian et al., 2024) have conducted thorough empirical evaluations of how preference alignment algorithms enhance LM-based TTS, and Seed-TTS (Anastassiou et al., 2024) also have adopted RLHF during the post-training stage of LM-based TTS. However, current approaches are limited to utterance-level preference optimization and overlook the potential of fine-grained alignment.
3 Methodology
3.1 Background
Neural codec based zero-shot TTS comprises two components: a neural codec model and a language model. The codec model tokenizes the speech signal into discrete acoustic tokens $a\in\mathbb{R}^{n\times l}$, where $n$ is the number of quantizer layers in the codec model and $l$ is the temporal length. The tokenized acoustic tokens can later be reconstructed into speech signals by the codec decoder. Given a text transcript $t$, the language model predicts the acoustic tokens using a next-token prediction approach. A widely used codec model employs a single quantizer, resulting in the extracted acoustic token being a single sequence along the temporal dimension. This setup simplifies the process of directly concat $t$ and $a$ and computing their joint probabilities in an auto-regressive manner:
$p(x)=\prod_{i=1}^{l}p\left(a_{l}\mid a_{1},\ldots,a_{l-1},t\right),$ (1)
it can also be regarded as a speech continuation task, conditioned on both the text and prompt speech during inference.
Figure 2: The preference optimization process of our proposed FPO. The language model is optimized using a selective training loss strategy, which focuses loss on the problematic segments while avoiding well-learned tokens.
RLHF with Preference Data. The language model is prompted with prompts $x$ to generate pairs of speech samples ($y_{1}$,$y_{2}$). Humans then evaluate these pairs and express preferences for one sample, denoted as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ represent the preferred and dispreferred samples respectively. The probability of preference data can be modeled using a reward model, which is trained via maximum likelihood estimation. During the reinforcement learning phase, the reward model provides feedback to the language model, which is then optimized by maximizing the rewards.
3.2 Preference Data Collection
The selection of preference data plays a critical role in RLHF. We propose a fine-grained sampling-annotation pipeline that begins with annotating pairwise preference data. Specifically, we first sample speech prompts from an unseen speaker pool for each text transcript and use them as input to the TTS system. To promote more diverse zero-shot TTS generation, we adjust sampling hyperparameters during inference and perform multiple inference runs for each sample, constructing a generated sample dataset $\{y_{1}^{b},y_{2}^{b},\ldots,y_{k}^{b}\}$, where $b$ is the $b$-th inference process.
After the sampling process, human annotators evaluate the generated samples from each inference process, identifying preferred and dispreferred ones. To mitigate the high human resource demands for annotations, we draw inspiration from previous studies (Lin et al., 2024a) and use pre-trained speech perceptual systems to simulate human feedback. We propose a comprehensive scoring method that combines multiple metrics—including intelligibility, quality, similarity, and duration—to determine the final score of each sample. This scoring system enables the identification of preferred samples based on their final scores $s(y)$:
$s(y)=\lambda_{w}\cdot w(y)^{p}+\lambda_{m}\cdot m(y)^{p}+\lambda_{c}\cdot c(y)^{p}+\lambda_{d}\cdot\text{Dur}(y)^{p},$ (2)
where $w(\cdot)$ and $m(\cdot)$ represent the intelligibility evaluation metric and quality evaluation metric, respectively. $c(y)$ measures the speaker similarity and $\text{Dur}(y)$ measures the duration difference between the generated samples and ground truth samples. The weights $\lambda_{*}$ normalize each metric to ensure they operate at the same amplitude and scale the range of $s(y)$ at [0,1]. Meanwhile, $p$ is used to increase sensitivity to metrics with small changes. These metrics are computed using pre-trained automatic systems, simulating various aspects of human evaluation. Therefore, the selection of preference data can be described as follows:
$(y_{l},y_{w})\in\Big{\{}y_{l}=\min$$\{s(y)\},\,y_{w}=\max\{s(y)\},$
$y_{w}-y_{l}>\tau\Big{\}},$ (3)
where $\tau$ is introduced to constrain sufficient differences between the preferred and dispreferred samples. We set $\tau=0.3$ to enhance the reliability of preference data for training.
3.3 Fine-grained Annotation
To achieve fine-grained alignment of TTS systems, we first thoroughly investigate the characteristics of segmental errors and categorize them into two types: temporal modeling errors and semantic-phonetic alignment errors, details can be found in Appendix A.1. As illustrated in Figure 1, temporal modeling errors arise from the model’s inability to accurately capture the temporal dynamics of speech, resulting in issues such as mispronunciations, abnormal silences, and unnatural prosody. For mispronunciations and abnormal silences, we utilize Whisper-timestamped [^1] https://github.com/linto-ai/whisper-timestamped and Voice Activity Detection systems to annotate the timestamps of these errors. For unnatural prosody, we incorporate the capabilities of text LLMs and professional human annotators to precisely identify affected segments (details provided in the Appendix A.2). This allows us to obtain timestamps for problematic segments and annotate fine-grained preference labels for temporal modeling errors at the token level.
Semantic-phonetic alignment errors, on the other hand, arise from the mapping between phonetic input and acoustic output. These errors often manifest as repetitions or truncations, where the TTS system either over-relies on its previous outputs or fails to encode essential phonetic information. Annotating these error segments relies on automatic speech recognition systems and ground truth transcripts with forced alignment. However, unlike temporal modeling errors, which affect only specific segments, repetition and truncation errors often cause cascading issues after the initial error segment. Therefore, we apply different annotation strategies for each error type. For temporal modeling errors, only the problematic segments are annotated as error segments. In contrast, for semantic-phonetic alignment errors, all segments following the onset of the issue are annotated as error segments. The token-level preference annotation can be described as an indicator function:
$I\left(y^{i}\right)=\begin{cases}1&\text{ if }y^{i}\text{ in the error segment}\\ 0&\text{ otherwise }\end{cases},$ (4)
where $i$ represents the $i$-th token of an utterance preference data.
3.4 Preference Optimization
The preference optimization process of our proposed FPO is illustrated in Figure 2. A typical preference optimization process of RL with a reward model $r_{\phi}(x,y)$ can be formulated as
$\max_{\pi_{\theta}}\mathbb{E}\left[r_{\phi}(x,y)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\mathrm{ref}}(y\mid x)\right],$ (5)
where $\beta$ is a parameter controlling the deviation between the TTS model $\pi_{\theta}$ and the base reference model $\pi_{\text{ref}}$. However, estimating a reward model is computationally expensive and introduces additional complexity during training. To address this, we adopt the DPO (Rafailov et al., 2024) approach to align the TTS system, eliminating the need for a reward model.
Following previous studies (Korbak et al., 2022; Go et al., 2023), the optimal solution to the KL-constrained reward maximization objective in Eq.(5) can be described as follows:
$\pi_{r}(y\mid x)=\frac{1}{Z(x)}\pi_{\text{ref }}(y\mid x)\exp\left(\frac{1}{\beta}r(x,y)\right).$ (6)
where $r(x,y)$ is the reward function widely defined as $r(x,y)=r_{\phi}(x,y)-\beta(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x))$, while $r(x,y)$ in optimization can be expressed in terms of the corresponding optimal policy, the reference policy, and the unknown partition function $Z(\cdot)$. This reward function is formulated as
$r(x,y)=\beta\log\frac{\pi_{r}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}+\beta\log Z(x).$ (7)
Since the Bradley-Terry model depends only on the difference in rewards between two completions, the partition function $Z(\cdot)$ can be canceled, allowing the human preference probability to be expressed solely in terms of the optimal TTS model and the corresponding reference model. Finally, to achieve fine-grained optimization, we introduce the token function from Eq. (4) to compute selective training loss for both preferred and dispreferred samples:
$\mathcal{L}_{\text{FPO}}=-\mathbb{E}\Bigg{[}$ $\sum_{i}I(y^{i})\cdot\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$
$-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\Big{)}\Bigg{]}.$ (8)
This function allows us to fit a token-level implicit reward to optimize the TTS model using preference data while avoiding computational waste on well-trained tokens.
4 Experimental Setup
Table 1: Objective results on WER/CER, UTMOS, SECS, and bad case ratio with different languages. For WER/CER and bad case ratio, lower values indicate better performance, while for UTMOS and SECS, higher values are better. The best results in each category are highlighted in bold.
[…]
CN EN CN EN CN EN CN EN
GT 3.06 2.17 - 3.92 4.41 -
CosyVoice 8.26+0.0% 6.97+0.0% 0.65 0.68 3.74 3.98 31% 24%
SDPO 7.76-6.1% 6.54-6.2% 0.68 0.70 3.85 4.04 17% 13%
UNO 5.36-35.1% 4.79-31.3% 0.70 0.73 3.94 4.03 13% 12%
RIO 4.81-41.8% 4.45-36.2% 0.72 0.71 4.02 4.11 11% 9%
FPO 3.92-52.5% 3.15-54.8% 0.66 0.68 3.83 4.01 8% 5%
4.1 Dataset
We chose CosyVoice (Du et al., 2024) as our backbone model, which was trained on a large-scale dataset comprising approximately 170,000 hours. For RLHF optimization, we use the training set of LibriTTS (Zen et al., 2019) and AISHELL-3 (Shi et al., 2020) corpora, already included in the base model’s training. This setup eliminates the influence of introducing additional high-quality data. Following prior research (Hu et al., 2024), we sample a pool of speech prompts and a separate pool of target transcripts to generate preference samples using pre-trained CosyVoice model for RLHF optimization. Using the proposed fine-grained sampling-annotation pipeline, we select pairs of samples with significant differences as training data for RLHF, resulting in 1,000 preference training utterances for preference optimization.
4.2 Baseline Systems
In addition to using CosyVoice as the base model, we reproduce the following powerful preference optimization approaches based on CosyVoice as baseline systems for a comprehensive comparison:
  • Speech-DPO (Rafailov et al., 2024): We adapt the popular DPO algorithm to align the base model with human feedback, where preference data is optimized at the utterance level, similar to the method described in (Zhang et al., 2024), and denote it as SDPO.
  • UNO (Chen et al., 2024a): A recent study proposes a RLHF framework tailored for TTS optimization, specifically considering the uncertainty arising from the inherent variability in subjective human speech perception and evaluation.
  • RIO (Hu et al., 2024): It proposes a reverse inference optimization method based on the Bayesian principle to select exemplars from speech samples generated by the TTS system, enabling preference optimization to enhance the robustness of the TTS system.
4.3 Evaluation Metrics
Subjective Metrics: We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples. We invited 15 participants to listen to the audio samples and rate the naturalness on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural. Furthermore, we conduct an ABX test where participants listen to two generated speech samples from different models but based on the same input and then choose the sample that sounds more natural. If the samples are too close to distinguish, they are instructed to indicate a tie.
Objective Metrics: For objective evaluation, we use the speaker embedding cosine similarity (SECS), calculated using pre-trained automatic speaker verification models [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, to evaluate speaker similarity. Word error rate (WER) and character error rate (CER), transcribed by an automatic speech recognition model [^3] https://huggingface.co/openai/whisper-medium, are employed to evaluate robustness for Mandarin and English datasets, respectively. Additionally, speech quality is measured using a predicted mean opinion score (UTMOS), computed by a neural network model [^4] https://github.com/sarulab-speech/UTMOS22. Meanwhile, following previous studies (Hu et al., 2024), we introduce the bad case ratio to evaluate the robustness of different baseline systems, where samples with UTMOS below 3 or WER exceeding 20% are classified as bad cases.
5 Experimental Results
5.1 Objective Evaluation
We first compare our proposed FPO with baseline systems using objective metrics, results are shown in Table 1. For WER and CER results, FPO achieves significant improvements in speech intelligibility that reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) compared to the base model CosyVoice. This outperforms all baselines, including RIO (--41.8% CER and --36.2% WER) and UNO (--35.1% CER and --31.3% WER), demonstrating FPO’s superior ability to mitigate segmental errors such as mispronunciations and repetitions.
On the other hand, FPO demonstrates remarkable robustness in terms of the bad case ratio, reducing it to 8% (CN) and 5% (EN), outperforming all baselines by a large margin. This highlights FPO’s ability to effectively minimize segmental errors by optimizing problematic segments. These results demonstrate the efficacy of fine-grained preference optimization in addressing segmental errors and enhancing the overall reliability of TTS systems.
For metrics such as SECS and UTMOS, our proposed FPO achieves only slightly higher results than the base model, while RIO and UNO achieve better performance than FPO. This is because FPO focuses on addressing localized errors rather than global attribute alignment, such as timbre or overall quality. Since the base model already demonstrates strong speaker similarity and speech quality, improvements from baseline systems are marginal, serving as a minor enhancement. We believe this trade-off is justified, as FPO prioritizes resolving segmental errors, emphasizing improvements in intelligibility and robustness over enhancements in global attributes.
Figure 3: Violin plots of the subjective NMOS for speech generated between baseline systems and FPO.
5.2 Subjective Evaluation
To further evaluate the effectiveness of FPO, we conducted subjective evaluations on the naturalness of the generated samples. As shown in Figure 3, the violin plots illustrate the NMOS distribution, comparing the performance of FPO with baseline systems. The distribution range of CosyVoice is wider than that of other systems, with elongated density tails indicating the presence of notable outliers. This suggests instability in the naturalness performance of the generated samples. In contrast, the distribution range of FPO is more compact than CosyVoice and SDPO, demonstrating improved robustness after preference optimization. Although UNO and RIO also achieve relatively compact NMOS distributions, FPO outperforms UNO in interquartile range and shows a better overall NMOS distribution compared to RIO. A comparative analysis of the violin plots reveals significant differences between FPO and the baseline systems, particularly in median values and distribution patterns. These results highlight the superior performance of FPO in achieving more natural speech generation.
Figure 4: Results for ABX preference test, while "Base" denote the original CosyVoice model.
Figure 5: Results of FPO and baseline systems using different sizes of training data.
Apart from NMOS, we also conduct ABX preference tests to verify the performance gains of our proposed FPO compared to baseline systems. As shown in Figure 4, compared to the base model, FPO is preferred in 54.3% of cases and ties in 24.1%, significantly outperforming the base model in preference. Notably, FPO also surpasses SDPO by a considerable margin (44.2% vs. 28.7%, with 27.1% ties), demonstrating the effectiveness of fine-grained optimization compared to full-sequence optimization. The ABX test results further highlight FPO’s advantages.
On the other hand, FPO achieves a higher preference score than UNO and RIO in the ABX test. While UNO and RIO focus on enhancing overall attributes such as timbre and quality, FPO specifically targets problematic segments, reducing abrupt errors and improving local consistency. The higher preference scores in ABX tests suggest that fine-grained optimization of local segments better aligns with human perceptual priorities.
These results indicate that while utterance-level optimization methods effectively align the base model to some extent, FPO directly addresses critical pain points in the human listening experience. This leads to higher practical utility, even when objective metric scores are comparable. The strong alignment between subjective and objective evaluations further validates the effectiveness of the proposed FPO.
5.3 Analysis on Data Efficient
To further evaluate the data efficiency of our proposed FPO, we compare its performance with baseline systems using small-scale Mandarin training data, with CER and bad case ratio as evaluation metrics. As shown in Figure 5, FPO achieves a similar CER performance as the RIO system while using only 200 utterances, compared to the 800 utterances required by the baselines. FPO is up to four times faster than the baseline systems. Additionally, FPO outperforms all baselines in bad case ratio across different data sizes and achieves a speedup of approximately 4x compared to UNO.
Furthermore, we scale up the preference data size to compare the data scalability between the SDPO approach and our proposed FPO. As shown in Table 2, with just 1,000 training utterances, FPO reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) for CN and EN, respectively. This outperforms even SDPO trained on 5,000 utterances (CER: 4.14 vs. 3.11; WER: 3.85 vs. 2.56), highlighting FPO’s ability to maximize efficiency with limited data. Notably, FPO reduces the bad case ratio to 8% (CN) and 5% (EN) with only 1,000 samples, whereas SDPO requires 5,000 samples to achieve 9% (CN) and 8% (EN). These results further demonstrate FPO’s effectiveness in mitigating critical errors while maintaining high data efficiency.
Table 2: Comparison results between FPO and SDPO on several larger datasets. "Num" represents the number of utterances used for training.
Num (#) CER/WER Bad case ratio
CN EN CN EN
CosyVoice - 8.26 6.97 31% 24%
SDPO 1000 7.76 6.54 17% 13%
2000 6.06 5.79 14% 12%
5000 4.14 3.85 9% 8%
FPO 1000 3.92 3.15 8% 5%
2000 3.79 2.84 5% 4%
5000 3.11 2.56 3% 2%
FPO’s advantage stems from its selective training loss strategy, which focuses optimization on error segments identified through fine-grained annotations. By avoiding unnecessary updates to well-learned tokens, FPO reduces computational redundancy and accelerates convergence. In contrast, utterance-level methods like SDPO apply uniform optimization across entire sequences, diluting the impact of critical segments and requiring larger datasets to achieve comparable improvements.
6 Conclusion
In this study, we propose FPO, a novel RLHF framework designed to address localized errors in zero-shot TTS systems. By categorizing generation errors into systemic and segmental types and introducing token-level selective loss optimization, FPO aligns problematic segments using fine-grained preference data while avoiding computational waste on well-learned tokens. Experimental results demonstrate that FPO significantly improves speech intelligibility and robustness, reducing CER by 52.5% and WER by 54.8% compared to the base model. Additionally, FPO lowers the bad case ratio to 8% (CN) and 5% (EN), outperforming utterance-level preference optimization approaches. Subjective evaluations further confirm FPO’s superiority in naturalness and alignment with human perceptual priorities. Notably, FPO achieves comparable performance with 4× fewer training samples than baseline systems, highlighting its remarkable data efficiency.
Limitations
One of the main limitations of this study lies in the construction of the preference training dataset. The proposed framework relies on fine-grained manual annotation to ensure the quality and consistency. This process is both time-consuming and labor-intensive, which may limit the scalability of the approach to larger or more diverse datasets. Although our proposed approach is data efficient and need less annotation data compared with previous works, the fine-grained annotation at each sample is more difficult than utterance level annotation. In future work, we plan to explore strategies to alleviate this limitation. For example, leveraging weakly supervised, semi-supervised, or self-supervised learning techniques may help reduce the dependence on manually annotated data.
Ethics Statement
This work does not raise any ethical concerns. All data and pre-trained models used in this study are publicly available and comply with the following licenses: Creative Commons BY 4.0, Creative Commons CC0, Creative Commons BY-NC-ND 4.0, and Creative Commons BY-SA 4.0.
Appendix A Appendix
A.1 Segmental Errors
Temporal modeling errors stem from the model’s limitations in capturing the temporal dynamics of speech, leading to issues such as mispronunciation, where abrupt or unstable pronunciation changes occur due to insufficient modeling of pitch continuity or variation trends. Another common problem is abnormal silence, where the model erroneously inserts silent frames or extends silent segments, disrupting the naturalness of generated speech. Additionally, the generated speech may exhibit unnatural prosody (i.g. unnatural pause), as the model may fail to adequately capture local rhythmic patterns, including stress, duration, and speaking rate, resulting in speech that deviates from natural prosody.
Semantic-phonetic alignment errors arise from mapping between the phonetic input and the acoustic output. These errors often manifest as repetition, where the model over-relies on its previous outputs and redundantly generates the same words or phonemes. Alternatively, truncation errors may occur when the language model fails to encode key phonetic information from the input or neglects certain critical elements, leading to synthesized speech that lacks completeness and accuracy.
A.2 Prosody Annotation
For prosody issues, we primarily focus on abnormal pauses and annotate fine-grained prosody errors through a three-step process. First, we use an automatic speech recognition system to transcribe the speech into text. Next, we employ text LLMs to determine ideal pauses in the text and insert break tags. Finally, professional annotators evaluate the speech samples based on the break tags, identifying preferred and dispreferred prosody samples. Fine-grained timestamps are then obtained using force-alignment tools based on the text and break tags.

### v2

@@
-Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech
+FPO: Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech
 Abstract
-Integrating human feedback to align text-to-speech (TTS) system outputs with human preferences has proven to be an effective approach for enhancing the robustness of language model-based TTS systems.
-Current approaches primarily focus on using preference data annotated at the utterance level.
-However, frequent issues that affect the listening experience often only arise in specific segments of audio samples, while other segments are well-generated.
+Integrating reinforcement learning to align generated speech with human preferences has proven effective in improving the robustness of language model-based text-to-speech (TTS) systems.
+Current approaches primarily rely on preference data annotated at the utterance level.
+However, frequent issues affecting the listening experience, such as unnatural prosody, mispronunciations, or semantic inconsistencies, often arise only in specific segments of audio, while other segments may be well-generated and require no correction.
+This mismatch between coarse-grained supervision and fine-grained quality variation limits the effectiveness of preference-based optimization.
 In this study, we propose a fine-grained preference optimization approach (FPO) to enhance the robustness of TTS systems.
-FPO focuses on addressing localized issues in generated samples rather than uniformly optimizing the entire utterance.
-Specifically, we first analyze the types of issues in generated samples, categorize them into two groups, and propose a selective training loss strategy to optimize preferences based on fine-grained labels for each issue type.
-Experimental results show that FPO enhances the robustness of zero-shot TTS systems by effectively addressing local issues, significantly reducing the bad case ratio, and improving intelligibility.
-Furthermore, FPO exhibits superior data efficiency compared with baseline systems, achieving similar performance with fewer training samples.
-Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech
-Jixun Yao1, Yuguang Yang2, Yu Pan2, Yuan Feng2, Ziqian Ning1, Jianhao Ye2, Hongbin Zhou2, Lei Xie1 1Northwestern Polytechnical University 2Everest Team, Ximalaya
-1 Introduction
-Large language models (LLMs) demonstrate remarkable in-context learning abilities in natural language processing (NLP) tasks by scaling both model parameters and dataset size (Brown et al., 2020; Touvron et al., 2023; Grattafiori et al., 2024).
-Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017) is crucial for calibrating these models, aligning them with human preferences, and minimizing the risk of generating harmful or toxic content.
-This technique is essential for the success of systems like ChatGPT (Achiam et al., 2023) and has recently been extended to tasks in other modalities.
-Speech, a key modality of generative artificial intelligence, is experiencing rapid advancements inspired by developments in text-based LLMs.
-Recent text-to-speech (TTS) studies (Wang et al., 2023; Chen et al., 2024b; Du et al., 2024) tokenize speech signals into discrete units and employ language model to generate speech tokens in a next-token prediction manner based on phoneme inputs and prompt speech recordings.
-By leveraging in-context learning capabilities, these TTS systems can produce high-quality, personalized speech in a zero-shot manner.
-Similar to the use of RLHF in calibrating text LLMs, there is growing interest in integrating RLHF into TTS systems (Zhang et al., 2024; Chen et al., 2024a; Hu et al., 2024; Tian et al., 2024).
-These approaches effectively enhance the robustness and zero-shot capabilities of TTS systems by incorporating human evaluation.
-Despite significant advancements in integrating RLHF into TTS systems, we believe two primary challenges remain.
+FPO shifts the optimization paradigm from global utterance-level tuning to targeted local refinement, focusing on addressing localized issues in generated samples rather than uniformly optimizing the entire utterance.
+We begin by analyzing the types of common generation issues and categorizing them into temporal modeling errors and semantic-phonetic alignment errors, which frequently degrade intelligibility and naturalness.
+To tackle these problems, we introduce a selective training loss strategy that leverages fine-grained labels for each issue type, allowing the model to focus on learning signals where they are most needed.
+Experimental results demonstrate that FPO substantially improves the robustness of zero-shot TTS systems by effectively correcting problematic regions in the output.
+This leads to a significant reduction in the bad case ratio, improved intelligibility, and overall perceptual quality.
+Moreover, FPO exhibits strong data efficiency, achieving comparable or superior performance to baseline methods while requiring notably fewer training samples.
+Audio samples can be found at https://yaoxunji.github.io/fpo/.
+I Introduction
+Large language models (LLMs) demonstrate remarkable in-context learning abilities in natural language processing (NLP) tasks through scaling model parameters and dataset sizes [brown2020language, touvron2023llama, grattafiori2024llama3].
+This scaling has enabled them to generalize across a wide range of tasks with minimal supervision, often achieving near-human or even superhuman performance in many domains.
+However, their vast and diverse training data encompasses varied goals, priorities, and content, not all of which are desirable to imitate.
+Selecting desired responses and behaviors from the model’s broad knowledge and capabilities is crucial for developing safe, high-performing, and controllable AI systems [christiano2017deep].
+A key solution involves aligning LLMs with human preferences through reinforcement learning (RL), effectively calibrating outputs to reduce harmful content risks.
+This approach underpins widely-used systems like ChatGPT [achiam2023gpt] and has recently expanded beyond pure language generation to multimodal tasks, including vision-language understanding [kelly2024visiongpt], code generation [zheng2024opencodeinterpreter], and speech processing [yao2025gense].
+Speech, a key modality in generative artificial intelligence, is experiencing rapid advancements inspired by developments in text-based LLMs.
+Recent text-to-speech (TTS) studies [wang2023valle, chen2024vall2, du2024cosyvoice] tokenize speech signals into discrete units via neural codec models, aligning them with phoneme or text inputs.
+These tokens are then processed by a language model to predict subsequent tokens in a next-token prediction manner using phoneme inputs and speech prompts.
+By scaling up model parameters and training data size, the in-context learning capabilities also emerge in TTS language models, enabling zero-shot generation of high-quality, personalized speech.
+Mirroring RL’s role in refining text LLMs, emerging interest focuses on applying RL to TTS systems [zhang2024speechalign, chen2024uno, hu2024rio, tian2024tx_rl].
+Typical approaches introduce direct preference optimization (DPO) [rafailov2024dpo] or Kahneman-Tversky Optimization (KTO) [ethayarajh2024kto] to align TTS systems, while the key challenge is to collect desirable and relevant preference datasets.
+By annotating generated speech samples using professional annotators at the utterance-level, these approaches effectively enhance the speech quality and zero-shot capabilities of TTS systems by post-training in an RL manner.
+Despite significant advancements in integrating RL into TTS systems, we believe two primary challenges remain.
 The first challenge is efficiently calibrating pre-trained TTS systems with limited preference data.
-RLHF typically requires preference data evaluated by human annotators, which is then used for training to align TTS systems with human preferences.
+RL in TTS studies typically requires preference data evaluated by human annotators, which is then used to align TTS systems with human preferences.
 However, obtaining large-scale human-annotated speech preference data is prohibitively expensive.
@@
 The second challenge is achieving fine-grained alignment of specific segments in generated samples with human feedback.
-In recent years, issues such as poor speaker similarity and degraded audio quality, referred to as systemic errors, have been largely mitigated in TTS systems through advanced technology.
-However, more frequent issues that impact the listening experience often only occur in specific segments of the audio samples, which we call segmental errors.
-These include issues like mispronunciation of certain words, abnormal pauses, or repetition, etc.
-However, current speech RLHF approaches typically annotate preference data at the utterance level, which fails to account for fine-grained issues.
-Additionally, utterance-level preference optimization calculates the loss for the entire sequence, even though many tokens are already well-trained, leading to inefficiencies (Lin et al., 2024b).
-We believe that optimizing problematic segments in a fine-grained manner would be both more effective and more data-efficient.
-To address the above challenges, we propose a novel fine-grained preference optimization (FPO) framework—an RLHF-based approach aimed at addressing segmental errors in generated samples and improving the robustness of zero-shot TTS systems.
-FPO employs a fine-grained sampling-annotating pipeline, where representative audio samples are first generated multiple times by TTS systems.
-Preference data are then annotated from human listening perspectives, labeling each sample as preferred or dispreferred at the utterance level.
-For each paired preference data, we annotate the problematic segment of the local issue to obtain fine-grained token-level preference labels, enabling more precise optimization than traditional approaches.
-Furthermore, we analyze the types of issues across different problem segments from an evaluation perspective and propose a selective training loss strategy to optimize each type of issue.
-During training, FPO processes the entire sequence through the TTS models but computes loss exclusively for the problematic segments, which can effectively resolve fine-grained issues while enhancing training efficiency.
-Experimental results demonstrate that FPO significantly enhances the robustness of zero-shot TTS systems and exhibits superior data efficiency compared with baseline systems, improving the quality of generated speech across both subjective and objective metrics.
+While issues like poor speaker similarity and degraded audio quality—referred to as systemic errors—have largely been mitigated in TTS systems by scaling up training data, more frequent issues affecting the listening experience often occur in specific segments of the audio.
+We refer to these as segmental errors, which include mispronunciations, abnormal pauses, repetitions, and other disruptions.
+Current speech RL approaches typically annotate preference data at the utterance level, failing to capture these fine-grained issues.
+Moreover, using utterance-level preference data to optimize TTS systems needs to calculate losses of the entire sequence, even when many segments are already well-trained, leading to ineffective gradient updates and inefficient optimization [lin2024rho].
+We believe that a more fine-grained approach to optimizing problematic segments would be both more effective and data-efficient.
+To address the above challenges, we propose a novel fine-grained preference optimization (FPO) framework for TTS systems—an RL-based approach designed to address segmental errors in generated samples and enhance the robustness of zero-shot TTS systems.
+FPO employs a fine-grained sampling-annotation pipeline for preference data collection, which plays a crucial role in the RL-based approaches.
+Pre-trained TTS systems first generate representative audio samples multiple times as the candidate preference data, while the utterance-level preference and dispreference data are annotated from human listening perspectives using these candidate data.
+For each paired preference data, we further identify the problematic segment and annotate it with fine-grained token-level preference labels, enabling more precise optimization than traditional methods.
+Furthermore, we analyze the types of issues across different segments from an evaluation perspective and divide the two types of errors in the annotated dispreference data.
+A selective training loss strategy is proposed to optimize different types of issues during preference optimization.
+Specifically, FPO processes the entire sequence through the TTS models but computes loss exclusively for the problematic segments, which can effectively resolve fine-grained issues while enhancing training efficiency.
+Experimental results demonstrate that FPO significantly enhances the robustness of zero-shot TTS systems and exhibits superior data efficiency compared to baseline systems, improving the quality of generated speech across both subjective and objective metrics.
 The main contributions can be summarized as follows:
-  • We propose FPO, an RLHF-based approach designed to align problematic segments in generated samples, thereby enhancing the robustness of TTS systems.
-The key novelty of FPO lies in its preference optimization strategy, which focuses on optimizing specific problematic segments while avoiding unnecessary computation on well-learned tokens, resulting in a more data-efficient optimization process.
-  • By analyzing issues in generated samples, FPO fine-grained optimizes TTS systems on problematic segments using a selective training strategy.
-This approach introduces a new perspective for addressing fine-grained fixes in TTS preference optimization, focusing on resolving bad cases efficiently and effectively.
-  • Extensive experiments demonstrate the efficiency and effectiveness of FPO across both subjective and objective metrics.
-Notably, FPO significantly reduces the occurrence of bad cases compared to powerful baseline systems, effectively addressing the robustness challenges of zero-shot TTS systems.
-Audio samples can be found at https://yaoxunji.github.io/fpo.
-2 Related Work
-2.1 Neural codec based Zero-shot TTS
-Inspired by the success of LLMs in NLP, formatting the TTS task as a next-token prediction problem has gained significant popularity in recent years (Borsos et al., 2022; Zhang et al., 2023).
-VALL-E (Wang et al., 2023), as a pioneering work, introduced the first neural codec based TTS framework, leveraging large-scale speech data to develop in-context learning capabilities.
-It can synthesize high-quality personalized speech in a zero-shot manner using only a 3-second enrolled recording as an acoustic prompt.
-Building on this foundation, the typical neural codec-based TTS systems (Kharitonov et al., 2023; Xin et al., 2024; Peng et al., 2024) involve tokenizing speech signals into discrete tokens using a neural codec (Défossez et al., 2022; Zeghidour et al., 2021).
-A decoder-only language model is then trained to predict these discrete tokens alongside phoneme sequences.
-Finally, the predicted tokens are reconstructed into speech signals using the codec decoder.
-This approach has demonstrated promising scalability with large datasets and model sizes, enabling high-quality synthesized speech and powerful zero-shot capabilities for unseen speakers (Łajszczak et al., 2024; Du et al., 2024).
+  • We propose FPO, an RL-based approach designed to align problematic segments in generated samples, thereby enhancing the robustness of TTS systems.
+The key novelty of FPO lies in its fine-grained preference optimization, which focuses on optimizing specific problematic segments rather than entire utterances.
+This approach avoids unnecessary computation on well-learned tokens, resulting in a more data-efficient optimization process.
+  • By analyzing issues in generated samples, FPO optimizes TTS systems by aligning problematic segments through a selective training strategy.
+This approach offers a new perspective for addressing fine-grained fixes in TTS preference optimization, which resolves bad cases both efficiently and effectively.
+  • Both subjective and objective experiments demonstrate the efficiency and effectiveness of FPO.
+Notably, FPO significantly reduces the occurrence of bad cases compared to baseline systems and can achieve similar performance with fewer training samples.
+II Related Work
+II-A Neural codec based Zero-shot TTS
+Inspired by the success of LLMs in NLP, formatting the TTS task as a next-token prediction problem has gained significant popularity in recent years [borsos2022audiolm, zhang2023speechgpt].
+This formulation allows TTS systems to benefit from autoregressive modeling and in-context learning capabilities that have proven effective in text generation tasks.
+VALL-E [wang2023valle], as a pioneering work in this direction, introduced the first neural codec-based TTS framework, leveraging large-scale speech data to develop strong generalization and prompt-based adaptation abilities.
+It can synthesize high-quality, personalized speech in a zero-shot manner using only a 3-second enrolled recording as an acoustic prompt, without requiring speaker-specific fine-tuning or additional adaptation steps.
+Building on this foundation, a series of neural codec-based TTS systems [kharitonov2023speak, xin2024ralle, peng2024voicecraft] have been developed.
+These models typically involve a three-stage pipeline: (1) discretizing continuous speech waveforms into semantic and/or acoustic tokens using a neural audio codec [defossez2022high, zeghidour2021soundstream, zhang2023speechtokenizer, ji2024wavtokenizer], (2) training a decoder-only language model to predict these tokens conditioned on phoneme sequences and optional prompts, and (3) reconstructing the speech waveform from predicted tokens using the codec’s decoder.
+This discrete tokenization enables autoregressive modeling and makes the TTS task more analogous to language modeling.
+Another prevalent approach introduces self-supervised speech representation models [hsu2021hubert, chen2022wavlm, chung2021w2v] to extract semantic tokens that capture linguistic content in a speaker-agnostic manner.
+These semantic tokens serve as intermediate targets for language models, which eases the token prediction task compared to directly modeling raw acoustic tokens[defossez2022high].
+To reconstruct the waveform, many systems further employ diffusion-based vocoders or conditional decoders that synthesize speech from the generated semantic tokens, optionally guided by acoustic features or speaker embeddings.
+These neural codec-based frameworks have demonstrated strong scalability with increasing data and model size, enabling high-fidelity speech synthesis and robust zero-shot generalization to unseen speakers, speaking styles, and acoustic conditions [lajszczak2024base, du2024cosyvoice].
+Despite their impressive performance, these models still face challenges in terms of robustness, particularly under mismatched conditions, and often suffer from localized quality degradations that are not effectively addressed by global training objectives.
 Figure 1: Typical examples of segmental errors can be categorized into two types: temporal modeling errors and semantic-phonetic alignment errors.
 The problematic segments in the generated samples are highlighted using red boxes or symbols.
-2.2 Reinforcement Learning from Human Feedback
-Human feedback has been widely used in language models for NLP tasks (Ouyang et al., 2022), with RLHF technology playing a crucial role in aligning models with human preferences.
-Conventionally, preference optimization is performed by maximizing rewards computed from a reward model to align generated content.
-Recent approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), leverage closed-form loss functions that directly operate on preference data.
-This method eliminates the need for explicit reward model training while achieving alignment performance comparable to traditional RLHF.
-RLHF in TTS remains in its infancy compared to its advancements in NLP.
-SpeechAlign first introduced RLHF in TTS (Zhang et al., 2024), exploring multiple preference alignment methods using paired preference data.
-UNO (Chen et al., 2024a) and RIO (Hu et al., 2024) extended this by optimizing unpaired preference data: UNO accounts for annotation uncertainty in subjective evaluations, while RIO introduces a reverse preference data selection method based on Bayesian principles.
-Recent studies (Tian et al., 2024) have conducted thorough empirical evaluations of how preference alignment algorithms enhance LM-based TTS, and Seed-TTS (Anastassiou et al., 2024) also have adopted RLHF during the post-training stage of LM-based TTS.
-However, current approaches are limited to utterance-level preference optimization and overlook the potential of fine-grained alignment.
-3 Methodology
-3.1 Background
+II-B Learning from Human Feedback
+Human feedback has been widely utilized in LLMs for NLP tasks [ouyang2022training, stiennon2020learning], where RL techniques play a crucial role in aligning model behavior with human preferences.
+This alignment paradigm has been instrumental in developing AI agents that are helpful, harmless, and honest, paving the way for controllable language models.
+Traditionally, RL-based preference optimization follows a two-step pipeline.
+First, a reward model is trained on human-annotated preference data, where annotators compare model outputs and indicate which one is preferred.
+Then, the language model is fine-tuned using reinforcement learning, typically with Proximal Policy Optimization (PPO) [schulman2017proximal], to maximize the expected reward as estimated by the reward model.
+While effective, this approach is computationally intensive and sensitive to reward model quality and policy instability during RL fine-tuning.
+To address these limitations, recent approaches such as Direct Preference Optimization (DPO) [rafailov2024dpo] propose a more efficient alternative.
+DPO introduces a closed-form loss function that operates directly on preference pairs, eliminating the need for an explicit reward model or reinforcement learning loop.
+This formulation simplifies the training pipeline while achieving alignment performance comparable to traditional RLHF methods.
+Other variants [gu2025mask], such as Implicit Preference Optimization (IPO) [garg2025ipo] and Kahneman-Tversky Optimization (KTO) [ethayarajh2024kto], explore related directions to improve preference learning stability and generalization.
+Although these methods have been extensively applied to NLP tasks, their adoption in speech-related domains remains relatively limited.
+Recent works begin exploring learning from human preferences in speech synthesis and voice conversion, yet most rely on utterance-level binary labels, which can be insufficient to capture the fine-grained nature of perceptual quality issues in speech.
+Learning from human feedback in TTS remains in its infancy compared to its advancements in NLP.
+SpeechAlign first introduced RLHF in TTS [zhang2024speechalign], exploring multiple preference alignment methods using paired preference data.
+UNO [chen2024uno] and RIO [hu2024rio] extended this by optimizing unpaired preference data: UNO accounts for annotation uncertainty in subjective evaluations, while RIO introduces a reverse preference data selection method based on Bayesian principles.
+Recent studies [tian2024tx_rl] have conducted thorough empirical evaluations of how preference alignment algorithms enhance LM-based TTS, and Seed-TTS [anastassiou2024seedtts] has also adopted the RL method during the post-training stage of LM-based TTS.
+However, current RL-based speech synthesis approaches are limited to utterance-level preference optimization and overlook the potential of fine-grained alignment.
+While token-level alignment methods (e.g., Mask-DPO [gu2025maskdpo], Rho-1 [lin2024rho]) have been explored in NLP, we adapt and extend it to speech-specific issues such as temporal modeling and semantic-phonetic misalignment, which are absent in text-based tasks.
+III Methodology
+III-A Background
 Neural codec based zero-shot TTS comprises two components: a neural codec model and a language model.
@@
 Given a text transcript $t$, the language model predicts the acoustic tokens using a next-token prediction approach.
-A widely used codec model employs a single quantizer, resulting in the extracted acoustic token being a single sequence along the temporal dimension.
+A widely used codec model employs a single quantizer [ye2025llasa], resulting in the extracted acoustic token being a single sequence along the temporal dimension.
 This setup simplifies the process of directly concat $t$ and $a$ and computing their joint probabilities in an auto-regressive manner:
@@
 it can also be regarded as a speech continuation task, conditioned on both the text and prompt speech during inference.
-Figure 2: The preference optimization process of our proposed FPO.
-The language model is optimized using a selective training loss strategy, which focuses loss on the problematic segments while avoiding well-learned tokens.
-RLHF with Preference Data.
-The language model is prompted with prompts $x$ to generate pairs of speech samples ($y_{1}$,$y_{2}$).
+RL with Preference Data.
+The language model is first employed to generate paired speech samples ($y_{1}$$y_{2}$) with prompts $x$.
 Humans then evaluate these pairs and express preferences for one sample, denoted as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ represent the preferred and dispreferred samples respectively.
@@
 During the reinforcement learning phase, the reward model provides feedback to the language model, which is then optimized by maximizing the rewards.
-3.2 Preference Data Collection
-The selection of preference data plays a critical role in RLHF.
+III-B Preference Data Collection
+The selection of preference data plays a critical role in the RL process.
 We propose a fine-grained sampling-annotation pipeline that begins with annotating pairwise preference data.
 Specifically, we first sample speech prompts from an unseen speaker pool for each text transcript and use them as input to the TTS system.
-To promote more diverse zero-shot TTS generation, we adjust sampling hyperparameters during inference and perform multiple inference runs for each sample, constructing a generated sample dataset $\{y_{1}^{b},y_{2}^{b},\ldots,y_{k}^{b}\}$, where $b$ is the $b$-th inference process.
+To promote more diverse zero-shot TTS generation, we adjust sampling hyperparameters during inference and perform multiple inference times for each prompt, constructing a generated sample dataset $\{y_{1}^{b},y_{2}^{b},\ldots,y_{k}^{b}\}$, where $b$ is the $b$-th inference process.
 After the sampling process, human annotators evaluate the generated samples from each inference process, identifying preferred and dispreferred ones.
-To mitigate the high human resource demands for annotations, we draw inspiration from previous studies (Lin et al., 2024a) and use pre-trained speech perceptual systems to simulate human feedback.
+To mitigate the high human resource demands for annotations, we draw inspiration from previous studies [lin2024alignslm] and use pre-trained speech perceptual systems to simulate human feedback.
 We propose a comprehensive scoring method that combines multiple metrics—including intelligibility, quality, similarity, and duration—to determine the final score of each sample.
-This scoring system enables the identification of preferred samples based on their final scores $s(y)$:
+This scoring system can identify preferred samples based on their final scores $s(y)$:
 $s(y)=\lambda_{w}\cdot w(y)^{p}+\lambda_{m}\cdot m(y)^{p}+\lambda_{c}\cdot c(y)^{p}+\lambda_{d}\cdot\text{Dur}(y)^{p},$ (2)
@@
 where $\tau$ is introduced to constrain sufficient differences between the preferred and dispreferred samples. We set $\tau=0.3$ to enhance the reliability of preference data for training.
-3.3 Fine-grained Annotation
-To achieve fine-grained alignment of TTS systems, we first thoroughly investigate the characteristics of segmental errors and categorize them into two types: temporal modeling errors and semantic-phonetic alignment errors, details can be found in Appendix A.1.
-As illustrated in Figure 1, temporal modeling errors arise from the model’s inability to accurately capture the temporal dynamics of speech, resulting in issues such as mispronunciations, abnormal silences, and unnatural prosody.
-For mispronunciations and abnormal silences, we utilize Whisper-timestamped [^1] https://github.com/linto-ai/whisper-timestamped and Voice Activity Detection systems to annotate the timestamps of these errors.
-For unnatural prosody, we incorporate the capabilities of text LLMs and professional human annotators to precisely identify affected segments (details provided in the Appendix A.2).
+Figure 2: The preference optimization process of our proposed FPO.
+The language model is optimized using a selective training loss strategy, which focuses loss on the problematic segments while avoiding well-learned tokens.
+III-C Fine-grained Annotation
+To achieve fine-grained alignment of TTS systems, we first thoroughly investigate the characteristics of segmental errors and categorize them into two types: temporal modeling errors and semantic-phonetic alignment errors.
+As illustrated in Figure 1, temporal modeling errors stem from the model’s limitations in capturing the temporal dynamics of speech, leading to issues such as mispronunciation, where abrupt or unstable pronunciation changes occur due to insufficient modeling of pitch continuity or variation trends.
+Another common problem is abnormal silence, where the model erroneously inserts silent frames or extends silent segments, disrupting the naturalness of generated speech.
+Additionally, the generated speech may exhibit unnatural prosody (i.g.
+unnatural pause), as the model may fail to adequately capture local rhythmic patterns, including stress, duration, and speaking rate, resulting in speech that deviates from natural prosody.
+For mispronunciations and abnormal silences, we employ Whisper-timestamped [^1] https://github.com/linto-ai/whisper-timestamped and voice activity detection systems to annotate the timestamps of these errors.
+To address unnatural prosody, we leverage the capabilities of text LLMs and professional human annotators to precisely identify affected segments.
+Specifically, we focus on abnormal pauses and annotate fine-grained prosody errors through a three-step process.
+First, we use an automatic speech recognition system to transcribe the speech into text.
+Next, we employ text LLMs to determine optimal pause placements and insert break tags accordingly.
+Finally, professional annotators evaluate the speech samples using these break tags, identifying preferred and dispreferred prosody patterns.
+Fine-grained timestamps are then obtained using force-alignment tools based on the text and break tags.
 This allows us to obtain timestamps for problematic segments and annotate fine-grained preference labels for temporal modeling errors at the token level.
@@
 These errors often manifest as repetitions or truncations, where the TTS system either over-relies on its previous outputs or fails to encode essential phonetic information.
-Annotating these error segments relies on automatic speech recognition systems and ground truth transcripts with forced alignment.
+Alternatively, truncation errors may occur when the language model fails to encode key phonetic information from the input or neglects certain critical elements, leading to synthesized speech that lacks completeness and accuracy.
+Annotating these error segments relies on automatic speech recognition systems and ground truth transcripts with force alignment tools.
 However, unlike temporal modeling errors, which affect only specific segments, repetition and truncation errors often cause cascading issues after the initial error segment.
@@
 The token-level preference annotation can be described as an indicator function:
-$I\left(y^{i}\right)=\begin{cases}1&\text{ if }y^{i}\text{ in the error segment}\\ 0&\text{ otherwise }\end{cases},$ (4)
-where $i$ represents the $i$-th token of an utterance preference data.
-3.4 Preference Optimization
+[…]
+where $i$ represents the $i$-th token of preference data.
+In the fine-grained annotation process, preference data is first identified, followed by the annotation of fine-grained preference labels by human annotators or automated agents—steps not included in conventional annotation methods.
+While fine-grained annotation involves additional steps, it requires less data than the traditional approach.
+Although utterance-level annotation is faster for human annotators, it demands more data to achieve the same level of alignment as fine-grained preference optimization.
+Therefore, we believe the additional cost of fine-grained annotation is both more acceptable and more efficient than conventional utterance-level annotation.
+III-D Preference Optimization
 The preference optimization process of our proposed FPO is illustrated in Figure 2. A typical preference optimization process of RL with a reward model $r_{\phi}(x,y)$ can be formulated as
 $\max_{\pi_{\theta}}\mathbb{E}\left[r_{\phi}(x,y)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\mathrm{ref}}(y\mid x)\right],$ (5)
-where $\beta$ is a parameter controlling the deviation between the TTS model $\pi_{\theta}$ and the base reference model $\pi_{\text{ref}}$.
+where $\beta$ is a parameter controlling the deviation between the TTS model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$.
 However, estimating a reward model is computationally expensive and introduces additional complexity during training.
-To address this, we adopt the DPO (Rafailov et al., 2024) approach to align the TTS system, eliminating the need for a reward model.
-Following previous studies (Korbak et al., 2022; Go et al., 2023), the optimal solution to the KL-constrained reward maximization objective in Eq.(5) can be described as follows:
+To address this, we follow the DPO approach [rafailov2024dpo] to directly align the TTS system using preference data, eliminating the need for a reward model.
+Following previous studies [korbak2022reinforcement, go2023aligning], the optimal solution to the KL-constrained reward maximization objective in Eq.(5) can be described as follows:
 $\pi_{r}(y\mid x)=\frac{1}{Z(x)}\pi_{\text{ref }}(y\mid x)\exp\left(\frac{1}{\beta}r(x,y)\right).$ (6)
@@
 $r(x,y)=\beta\log\frac{\pi_{r}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}+\beta\log Z(x).$ (7)
-Since the Bradley-Terry model depends only on the difference in rewards between two completions, the partition function $Z(\cdot)$ can be canceled, allowing the human preference probability to be expressed solely in terms of the optimal TTS model and the corresponding reference model.
-Finally, to achieve fine-grained optimization, we introduce the token function from Eq.
+Since the Bradley-Terry model depends only on the difference in rewards between two completions, the partition function $Z(\cdot)$ can be canceled, allowing the human preference probability to be expressed solely in terms of the optimal TTS model and the reference model.
+Finally, to achieve fine-grained optimization, we introduce the indicator function from Eq.
 (4) to compute selective training loss for both preferred and dispreferred samples:
-$\mathcal{L}_{\text{FPO}}=-\mathbb{E}\Bigg{[}$ $\sum_{i}I(y^{i})\cdot\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$
-$-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\Big{)}\Bigg{]}.$ (8)
+$\mathcal{L}_{\text{FPO}}=-\mathbb{E}\Bigg[$ $\sum_{i}I(y^{i})\cdot\log\sigma\Big(\beta\log\frac{\pi_{\theta}(y_{w}^{i}|x)}{\pi_{\text{ref}}(y_{w}^{i}|x)}$
+$-\beta\log\frac{\pi_{\theta}(y_{l}^{i}|x)}{\pi_{\text{ref}}(y_{l}^{i}|x)}\Big)\Bigg].$ (8)
 This function allows us to fit a token-level implicit reward to optimize the TTS model using preference data while avoiding computational waste on well-trained tokens.
-4 Experimental Setup
-Table 1: Objective results on WER/CER, UTMOS, SECS, and bad case ratio with different languages.
-For WER/CER and bad case ratio, lower values indicate better performance, while for UTMOS and SECS, higher values are better.
-The best results in each category are highlighted in bold.
-[…]
-CN EN CN EN CN EN CN EN
-GT 3.06 2.17 - 3.92 4.41 -
-CosyVoice 8.26+0.0% 6.97+0.0% 0.65 0.68 3.74 3.98 31% 24%
-SDPO 7.76-6.1% 6.54-6.2% 0.68 0.70 3.85 4.04 17% 13%
-UNO 5.36-35.1% 4.79-31.3% 0.70 0.73 3.94 4.03 13% 12%
-RIO 4.81-41.8% 4.45-36.2% 0.72 0.71 4.02 4.11 11% 9%
-FPO 3.92-52.5% 3.15-54.8% 0.66 0.68 3.83 4.01 8% 5%
-4.1 Dataset
-We chose CosyVoice (Du et al., 2024) as our backbone model, which was trained on a large-scale dataset comprising approximately 170,000 hours.
-For RLHF optimization, we use the training set of LibriTTS (Zen et al., 2019) and AISHELL-3 (Shi et al., 2020) corpora, already included in the base model’s training.
+IV Experimental Setup
+TABLE I: Objective evaluation results on WER/CER, SECS, and bad case ratio with different languages.
+For WER/CER and bad case ratio, lower values indicate better performance, while for SECS, higher values are better.
+The results are averaged by five inference times and the best in each category are highlighted in bold.
+test-zh test-en test-hard UTMOS $\uparrow$ Bad Case (%) $\downarrow$ F0 Corr. $\uparrow$
+CER $\downarrow$ SECS $\uparrow$ WER $\downarrow$ SECS $\uparrow$ WER $\downarrow$ SECS $\uparrow$
+CosyVoice [du2024cosyvoice] 3.63-0.0% 0.723 4.29-0.0% 0.609 11.75-0.0% 0.709 3.38±0.11 21% 0.71
+   + SDPO [rafailov2024dpo] 2.69-25.9% 0.731 3.41-20.5% 0.645 9.84-16.3% 0.716 3.44±0.10 17% 0.70
+   + UNO [chen2024uno] 2.57-29.2% 0.735 3.06-28.7% 0.667 9.06-22.9% 0.722 3.67±0.11 15% 0.75
+   + RIO [hu2024rio] 2.21-39.1% 0.744 2.89-32.6% 0.681 9.12-22.4% 0.727 3.64±0.10 12% 0.73
+   + Mask-DPO [gu2025maskdpo] 2.61-28.1% 0.734 3.27-23.8% 0.652 9.52-19.0% 0.715 3.48±0.10 14% 0.65
+   + FPO (Ours) 1.83-49.6% 0.728 2.77-35.4% 0.614 7.98-32.1% 0.708 3.47±0.09 9% 0.78
+CosyVoice2 [du2024cosyvoice2] 1.45-0.0% 0.748 2.57-0.0% 0.652 8.08-0.0% 0.730 3.49±0.11 14% 0.75
+   + SDPO [rafailov2024dpo] 1.43-1.4% 0.745 2.71+5.4% 0.673 8.32+3.0% 0.728 3.59±0.08 15% 0.73
+   + UNO [chen2024uno] 1.37-5.5% 0.759 2.43-5.4% 0.669 7.78-3.7% 0.736 3.61±0.11 12% 0.75
+   + RIO [hu2024rio] 1.40-3.4% 0.751 2.47-3.9% 0.675 7.94-1.7% 0.737 3.73±0.12 9% 0.77
+   + Mask-DPO [gu2025maskdpo] 1.42-2.1% 0.754 2.51-2.3% 0.671 7.92-2.0% 0.733 3.64±0.10 13% 0.67
+   + FPO (Ours) 1.32-9.0% 0.747 2.24-12.8% 0.654 6.92-14.4% 0.731 3.56±0.10 8% 0.81
+Llasa [ye2025llasa] 1.89-0.0% 0.669 3.22-0.0% 0.572 12.13-0.0% 0.638 3.54±0.12 25% 0.73
+   + SDPO [rafailov2024dpo] 1.73-8.5% 0.685 2.83-12.1% 0.599 11.23-7.4% 0.633 3.67±0.11 22% 0.75
+   + UNO [chen2024uno] 1.64-13.2% 0.701 2.74-14.9% 0.623 10.84-10.6% 0.663 3.65±0.13 18% 0.74
+   + RIO [hu2024rio] 1.58-16.4% 0.709 2.79-13.4% 0.625 9.45-22.1% 0.671 3.75±0.11 15% 0.78
+   + Mask-DPO [gu2025maskdpo] 1.71-9.5% 0.694 2.84-11.8% 0.591 9.92-18.2% 0.637 3.64±0.10 19% 0.64
+   + FPO (Ours) 1.47-22.2% 0.672 2.46-23.6% 0.571 8.42-30.6% 0.640 3.58±0.12 11% 0.79
+IV-A Model Configuration
+We select the open-source CosyVoice [du2024cosyvoice], CosyVoice2 [du2024cosyvoice2] and Llasa [ye2025llasa] as our backbone models.
+These systems are trained on large-scale datasets and demonstrate impressive performance.
+CosyVoice series consists of a text encoder, a speech tokenizer, a speech language model and a conditional flow matching model.
+Specifically, the text encoder aligns the semantic spaces of text and speech tokens, while the speech tokenizer extracts semantic tokens.
+The speech language model then predicts speech tokens based on text inputs and an acoustic prompt.
+A conditional flow matching model converts the predicted speech tokens into a mel-spectrogram, which is finally reconstructed into speech signals by a pre-trained vocoder.
+In contrast, Llasa bypasses flow matching and the vocoder, directly reconstructing the predicted speech tokens into waveforms using a codec decoder.
+Our study focuses on optimizing the speech language model using fine-grained preference data.
+For model training, we use a learning rate of 2e-6 and a batch size of 64, optimized with the AdamW optimizer [loshchilov2017adamw].
+The model is trained for only two epochs.
+IV-B Dataset
+To construct the training data for preference optimization, we use the LibriTTS [zen2019libritts] and AISHELL-3 [shi2020aishell3] corpora, which are already included in the base models’ training.
 This setup eliminates the influence of introducing additional high-quality data.
-Following prior research (Hu et al., 2024), we sample a pool of speech prompts and a separate pool of target transcripts to generate preference samples using pre-trained CosyVoice model for RLHF optimization.
-Using the proposed fine-grained sampling-annotation pipeline, we select pairs of samples with significant differences as training data for RLHF, resulting in 1,000 preference training utterances for preference optimization.
-4.2 Baseline Systems
-In addition to using CosyVoice as the base model, we reproduce the following powerful preference optimization approaches based on CosyVoice as baseline systems for a comprehensive comparison:
-  • Speech-DPO (Rafailov et al., 2024): We adapt the popular DPO algorithm to align the base model with human feedback, where preference data is optimized at the utterance level, similar to the method described in (Zhang et al., 2024), and denote it as SDPO.
-  • UNO (Chen et al., 2024a): A recent study proposes a RLHF framework tailored for TTS optimization, specifically considering the uncertainty arising from the inherent variability in subjective human speech perception and evaluation.
-  • RIO (Hu et al., 2024): It proposes a reverse inference optimization method based on the Bayesian principle to select exemplars from speech samples generated by the TTS system, enabling preference optimization to enhance the robustness of the TTS system.
-4.3 Evaluation Metrics
+Following a prior study [hu2024rio], we sample a pool of speech prompts containing 100 individual speakers and a separate pool of target transcripts to generate preference samples using the pre-trained models.
+With the proposed fine-grained sampling–annotation pipeline, we then select sample pairs with substantial differences to build the training set, resulting in 1,000 utterances for preference optimization.
+For evaluation, we use the Seed-TTS eval dataset [anastassiou2024seedtts] to evaluate model performance after preference optimization.
+This test set includes samples extracted from publicly available English (EN) and Mandarin (ZH) corpora and is designed to measure performance across various objective metrics.
+Each sample in the English and Mandarin test sets consists of one reference utterance and one target utterance spoken by the same speaker.
+The optimized model generates speech for the target text using the reference speech as an audio prompt.
+IV-C Baseline Systems
+In addition to using the CosyVoice series as base models, we reproduce the following strong RL-based speech approaches based on CosyVoice to serve as powerful baseline systems for comprehensive comparison:
+  • Speech-DPO [rafailov2024dpo]: We adapt the popular DPO algorithm to align the base model with human feedback, optimizing preference data at the utterance level, similar to the approach described in [zhang2024speechalign].
+However, instead of using ground-truth and generated samples as pairs, we use both generated samples as the preference and the other as the dispreference.
+We refer to this baseline as SDPO.
+  • UNO [chen2024uno]: A recent study proposes an RL framework tailored for zero-shot TTS models, which integrates human feedback into the TTS learning objective and specifically considers the uncertainty arising from inherent variability in subjective human perception and evaluation.
+The preference optimization maximizes the utility of these samples in an uncertainty-aware manner.
+  • RIO [hu2024rio]: An RL-based preference optimization approach tailored to the zero-shot TTS system with neural codec modelling.
+It introduces a reverse inference optimization method based on the Bayesian principle to select exemplars from speech samples generated by the TTS system, which is further employed to generate the original prompt samples.
+By using the reverse inference method, it can effectively select preference data to enhance the capacity of an advanced zero-shot TTS system.
+  • Mask-DPO [gu2025maskdpo]: A fine-grained factuality alignment method based on DPO that incorporating sentence-level factuality as mask signals and only learns from factually correct sentences in the preferred samples and prevents the penalty on factual contents in the not preferred samples, which resolves the ambiguity in the preference learning.
+IV-D Evaluation Metrics
 Subjective Metrics: We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples.
 We invited 15 participants to listen to the audio samples and rate the naturalness on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural.
-Furthermore, we conduct an ABX test where participants listen to two generated speech samples from different models but based on the same input and then choose the sample that sounds more natural.
-If the samples are too close to distinguish, they are instructed to indicate a tie.
-Objective Metrics: For objective evaluation, we use the speaker embedding cosine similarity (SECS), calculated using pre-trained automatic speaker verification models [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, to evaluate speaker similarity.
-Word error rate (WER) and character error rate (CER), transcribed by an automatic speech recognition model [^3] https://huggingface.co/openai/whisper-medium, are employed to evaluate robustness for Mandarin and English datasets, respectively.
-Additionally, speech quality is measured using a predicted mean opinion score (UTMOS), computed by a neural network model [^4] https://github.com/sarulab-speech/UTMOS22.
-Meanwhile, following previous studies (Hu et al., 2024), we introduce the bad case ratio to evaluate the robustness of different baseline systems, where samples with UTMOS below 3 or WER exceeding 20% are classified as bad cases.
-5 Experimental Results
-5.1 Objective Evaluation
-We first compare our proposed FPO with baseline systems using objective metrics, results are shown in Table 1.
-For WER and CER results, FPO achieves significant improvements in speech intelligibility that reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) compared to the base model CosyVoice.
-This outperforms all baselines, including RIO (--41.8% CER and --36.2% WER) and UNO (--35.1% CER and --31.3% WER), demonstrating FPO’s superior ability to mitigate segmental errors such as mispronunciations and repetitions.
-On the other hand, FPO demonstrates remarkable robustness in terms of the bad case ratio, reducing it to 8% (CN) and 5% (EN), outperforming all baselines by a large margin.
-This highlights FPO’s ability to effectively minimize segmental errors by optimizing problematic segments.
-These results demonstrate the efficacy of fine-grained preference optimization in addressing segmental errors and enhancing the overall reliability of TTS systems.
-For metrics such as SECS and UTMOS, our proposed FPO achieves only slightly higher results than the base model, while RIO and UNO achieve better performance than FPO.
-This is because FPO focuses on addressing localized errors rather than global attribute alignment, such as timbre or overall quality.
-Since the base model already demonstrates strong speaker similarity and speech quality, improvements from baseline systems are marginal, serving as a minor enhancement.
-We believe this trade-off is justified, as FPO prioritizes resolving segmental errors, emphasizing improvements in intelligibility and robustness over enhancements in global attributes.
-Figure 3: Violin plots of the subjective NMOS for speech generated between baseline systems and FPO.
-5.2 Subjective Evaluation
-To further evaluate the effectiveness of FPO, we conducted subjective evaluations on the naturalness of the generated samples.
-As shown in Figure 3, the violin plots illustrate the NMOS distribution, comparing the performance of FPO with baseline systems.
-The distribution range of CosyVoice is wider than that of other systems, with elongated density tails indicating the presence of notable outliers.
-This suggests instability in the naturalness performance of the generated samples.
-In contrast, the distribution range of FPO is more compact than CosyVoice and SDPO, demonstrating improved robustness after preference optimization.
-Although UNO and RIO also achieve relatively compact NMOS distributions, FPO outperforms UNO in interquartile range and shows a better overall NMOS distribution compared to RIO.
-A comparative analysis of the violin plots reveals significant differences between FPO and the baseline systems, particularly in median values and distribution patterns.
-These results highlight the superior performance of FPO in achieving more natural speech generation.
-Figure 4: Results for ABX preference test, while "Base" denote the original CosyVoice model.
-Figure 5: Results of FPO and baseline systems using different sizes of training data.
-Apart from NMOS, we also conduct ABX preference tests to verify the performance gains of our proposed FPO compared to baseline systems.
-As shown in Figure 4, compared to the base model, FPO is preferred in 54.3% of cases and ties in 24.1%, significantly outperforming the base model in preference.
-Notably, FPO also surpasses SDPO by a considerable margin (44.2% vs.
-28.7%, with 27.1% ties), demonstrating the effectiveness of fine-grained optimization compared to full-sequence optimization.
-The ABX test results further highlight FPO’s advantages.
+We apply a t-test to assess the statistical significance between systems with 95% confidence intervals for the mean opinion scores to better reflect the variability of listener ratings.
+Furthermore, we conduct an ABX test in which participants listen to two generated speech samples from different models based on the same input and select the more natural-sounding sample.
+If the samples are too similar to distinguish, participants are instructed to indicate a tie.
+Objective Metrics: For objective evaluation, we use the speaker embedding cosine similarity (SECS), calculated using pre-trained WavLM-large fine-tuned automatic speaker verification models [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, to evaluate speaker similarity.
+For intelligibility evaluation, we employ the word error rate (WER) metric for English samples and the character error rate (CER) metric for Mandarin samples.
+English samples are transcribed using the Whisper-Large-v3 model [radford2022whisper] [^3] https://huggingface.co/openai/whisper-large-v3, while Mandarin samples are transcribed using the Paraformer model [gao2022paraformer] [^4] https://huggingface.co/funasr/paraformer-zh.
+Additionally, speech quality is further measured using a predicted mean opinion score (UTMOS), computed by a neural network model [^5] https://github.com/sarulab-speech/UTMOS22.
+Meanwhile, following previous studies [hu2024rio], we introduce the bad case ratio to evaluate the robustness of different baseline systems.
+Samples are classified as bad cases if their UTMOS is below 3, or if their WER exceeds 3% on the test-zh and test-en sets, or 6% on the test-hard set.
+Finally, we use the Pearson correlation (F0 Corr.) between the generated and ground-truth pitch sequences as a metric to evaluate prosody performance.
+V Experimental Results
+V-A Objective Evaluation
+We first compare our proposed FPO with baseline systems using objective metrics, results are shown in TableI.
+For WER and CER, FPO consistently achieves the most significant improvements in speech intelligibility across all backbones.
+With the CosyVoice model, FPO reduces CER by 49.6% (from 3.63 to 1.83) and WER by 35.4% (from 4.29 to 2.77), outperforming all baseline systems, including RIO (-39.1% CER, -32.6% WER) and Mask-DPO (-28.1% CER, -23.8% WER).
+Similarly, when built upon the stronger CosyVoice2 backbone, FPO further achieves the lowest CER (1.32) and WER (2.24 and 6.92 on the test-en and test-hard sets, respectively), surpassing other optimization methods such as UNO and SDPO.
+Notably, FPO also maintains desired SECS and UTMOS scores, indicating improved naturalness without compromising perceptual quality.Beyond CosyVoice-based models, we additionally evaluate FPO on Llasa, a more recent TTS backbone.
+FPO again demonstrates substantial intelligibility gains, reducing CER by 22.2% (from 1.89 to 1.47) and WER by 23.6% (from 3.22 to 2.46), while outperforming all other baselines, including Mask-DPO and RIO.
+In terms of robustness, FPO achieves the lowest bad case ratios across all settings—reducing them from 21% to 9% on CosyVoice, from 14% to 8% on CosyVoice2, and from 25% to 11% on Llasa, indicating a strong capability to suppress segmental degradation such as mispronunciations and unnatural repetitions.
+These consistent improvements across diverse architectures validate the generalizability and effectiveness of our fine-grained preference optimization strategy in enhancing both the reliability and perceptual consistency of speech synthesis systems.
+While FPO demonstrates only marginal improvements over the base model in metrics like SECS and UTMOS—with RIO and UNO outperforming it in these aspects—this discrepancy stems from the distinct optimization priorities of the approaches.
+FPO’s core design focuses on resolving localized segmental errors through token-level selective training, whereas baseline systems like RIO and UNO prioritize utterance-level alignment for global attributes such as speaker similarity and overall quality.
+The base model already achieves strong performance in speaker similarity and speech quality due to its large-scale pre-training, leaving limited room for improvement in these dimensions.
+Consequently, the gains from RIO and UNO—though statistically significant—represent incremental refinements rather than transformative advancements.
+In contrast, FPO’s targeted optimization reduces critical segmental errors by 40% in CER/WER and slashes the bad case ratio by a large margin, directly addressing the most frequent and perceptually salient issues in human evaluations.
+This trade-off aligns with practical priorities in real-world TTS deployment, where eliminating catastrophic local errors often outweighs minor improvements in global attributes.
+TABLE II: Subjective NMOS results for the generated speech between baseline systems and FPO with 95% confidence interval.
+CosyVoice CosyVoice2 Llasa
+Base 3.64±0.11 3.81±0.07 3.74±0.12
+SDPO 3.68±0.09 3.83±0.11 3.80±0.09
+Mask-DPO 3.60±0.12 3.79±0.09 3.71±0.11
+UNO 3.69±0.11 3.72±0.10 3.75±0.08
+RIO 3.76±0.08 3.87±0.08 3.82±0.09
+FPO 3.83±0.09 3.91±0.08 3.93±0.10
+Figure 3: Results of ABX preference test, while ”C1” and ”C2” denote CosyVoice and CosyVoice2 model, respectively. MDPO denotes Mask-DPO method.
+Figure 4: Results of FPO and baseline systems using different scales of training data.
+V-B Subjective Evaluation
+To further evaluate the effectiveness of FPO, we conduct subjective evaluations on the naturalness of the generated samples.
+We can find that our proposed FPO consistently achieves the highest NMOS scores among all baseline RLHF alignment approaches, as shown in Table II.
+While several baselines (e.g., RIO and SDPO) exhibit modest increases in naturalness, FPO provides the most substantial and stable enhancement, improving NMOS by +0.19, +0.10, and +0.19 points on CosyVoice, CosyVoice2, and Llasa, respectively.
+To further investigate the source of these perceptual gains, we combine the objective Pearson correlation results with the subjective evaluations and find that FPO outperforms all baseline systems on both metrics.
+This indicates that the improved naturalness is not merely due to better intelligibility or fewer segment-level errors.
+Instead, FPO more effectively aligns prosodic structure, directly contributing to the enhanced perceptual quality.
+These findings highlight FPO’s superior performance in generating more natural and consistent speech.
+In addition to NMOS, we conduct ABX preference tests to further validate the performance gains of our proposed FPO compared to baseline systems.
+The ABX tests include both CosyVoice and CosyVoice2 models, while all other baselines are trained based on CosyVoice2.
+As shown in Figure 3, FPO is preferred over CosyVoice in 57.1% of cases and ties in 21.4%.
+Compared to CosyVoice2, FPO achieves a 44.3% win rate, again significantly outperforming the base model in terms of listener preference.
+Notably, FPO also outperform SDPO by a substantial margin (40.8% vs.
+27.6%, with 31.6% ties), demonstrating the effectiveness of fine-grained optimization over full-sequence optimization.
+These ABX test results further emphasize the advantages of FPO in producing more preferred and natural speech.
 On the other hand, FPO achieves a higher preference score than UNO and RIO in the ABX test.
@@
 The higher preference scores in ABX tests suggest that fine-grained optimization of local segments better aligns with human perceptual priorities.
-These results indicate that while utterance-level optimization methods effectively align the base model to some extent, FPO directly addresses critical pain points in the human listening experience.
+These results indicate that while utterance-level optimization methods effectively align the base model to some extent, FPO directly addresses critical points in the human listening experience.
 This leads to higher practical utility, even when objective metric scores are comparable.
-The strong alignment between subjective and objective evaluations further validates the effectiveness of the proposed FPO.
-5.3 Analysis on Data Efficient
-To further evaluate the data efficiency of our proposed FPO, we compare its performance with baseline systems using small-scale Mandarin training data, with CER and bad case ratio as evaluation metrics.
-As shown in Figure 5, FPO achieves a similar CER performance as the RIO system while using only 200 utterances, compared to the 800 utterances required by the baselines.
-FPO is up to four times faster than the baseline systems.
-Additionally, FPO outperforms all baselines in bad case ratio across different data sizes and achieves a speedup of approximately 4x compared to UNO.
+The consistency between subjective and objective evaluations further validates the effectiveness of the proposed FPO.
+V-C Analysis on Data Efficient
+To further evaluate the data efficiency of our proposed FPO, we compare its performance against baseline systems using small-scale English training data, with intelligibility evaluated on the test-hard set using WER as metric and average bad case ratio across three test sets as another metric.
+As shown in Figure 4, FPO achieves comparable WER performance to the UNO system using only 200 training utterances, whereas the baselines require 600 utterances.
+This demonstrates that FPO is up to three times more data-efficient than baseline systems.
+Furthermore, FPO consistently outperforms all baselines in terms of bad case ratio across various data scales and achieves an approximate 4x speedup compared to UNO.
+TABLE III: Comparison results between FPO and SDPO on several larger-scale datasets. ”Num” represents the number of utterances used for training.
+Num (#) CER WER Bad case ratio
+CosyVoice2 - 1.45 5.33 14%
+ + SDPO 1000 1.43 5.52 15%
+2000 1.40 5.19 12%
+5000 1.31 3.94 7%
+ + FPO 1000 1.32 4.58 8%
+2000 1.29 3.76 5%
+5000 1.21 3.23 3%
 Furthermore, we scale up the preference data size to compare the data scalability between the SDPO approach and our proposed FPO.
-As shown in Table 2, with just 1,000 training utterances, FPO reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) for CN and EN, respectively.
-This outperforms even SDPO trained on 5,000 utterances (CER: 4.14 vs.
-3.11; WER: 3.85 vs.
-2.56), highlighting FPO’s ability to maximize efficiency with limited data.
-Notably, FPO reduces the bad case ratio to 8% (CN) and 5% (EN) with only 1,000 samples, whereas SDPO requires 5,000 samples to achieve 9% (CN) and 8% (EN).
+As shown in Table III, with just 1,000 training utterances, FPO reduces the CER of the original CosyVoice2 model from 1.45 to 1.32 and the WER from 5.33 to 4.58 for Mandarin and English evaluation datasets, respectively.
+This performance outperforms even SDPO trained on 5,000 utterances (CER: 1.32 vs.
+1.31; WER: 4.58 vs.
+3.94), highlighting FPO’s ability to maximize efficiency of preference optimization with limited data.
+Notably, FPO reduces the bad case ratio from 14% to 8% using only 1,000 samples, whereas SDPO requires 5,000 samples to achieve a similar performance of about 7%.
 These results further demonstrate FPO’s effectiveness in mitigating critical errors while maintaining high data efficiency.
-Table 2: Comparison results between FPO and SDPO on several larger datasets. "Num" represents the number of utterances used for training.
-Num (#) CER/WER Bad case ratio
-CN EN CN EN
-CosyVoice - 8.26 6.97 31% 24%
-SDPO 1000 7.76 6.54 17% 13%
-2000 6.06 5.79 14% 12%
-5000 4.14 3.85 9% 8%
-FPO 1000 3.92 3.15 8% 5%
-2000 3.79 2.84 5% 4%
-5000 3.11 2.56 3% 2%
-FPO’s advantage stems from its selective training loss strategy, which focuses optimization on error segments identified through fine-grained annotations.
-By avoiding unnecessary updates to well-learned tokens, FPO reduces computational redundancy and accelerates convergence.
-In contrast, utterance-level methods like SDPO apply uniform optimization across entire sequences, diluting the impact of critical segments and requiring larger datasets to achieve comparable improvements.
-6 Conclusion
-In this study, we propose FPO, a novel RLHF framework designed to address localized errors in zero-shot TTS systems.
-By categorizing generation errors into systemic and segmental types and introducing token-level selective loss optimization, FPO aligns problematic segments using fine-grained preference data while avoiding computational waste on well-learned tokens.
-Experimental results demonstrate that FPO significantly improves speech intelligibility and robustness, reducing CER by 52.5% and WER by 54.8% compared to the base model.
-Additionally, FPO lowers the bad case ratio to 8% (CN) and 5% (EN), outperforming utterance-level preference optimization approaches.
-Subjective evaluations further confirm FPO’s superiority in naturalness and alignment with human perceptual priorities.
-Notably, FPO achieves comparable performance with 4× fewer training samples than baseline systems, highlighting its remarkable data efficiency.
-Limitations
+FPO’s advantage stems from its selective training loss strategy, which precisely directs optimization toward error-prone segments identified through fine-grained annotations of temporal modeling and semantic-phonetic misalignments.
+By identifying problematic tokens and excluding well-learned regions from gradient updates, FPO minimizes computational redundancy while preventing over-optimization of already proficient segments.
+This targeted approach not only reduces noise in parameter updates but also accelerates convergence by prioritizing problematic segments.
+In contrast, utterance-level methods like SDPO apply uniform optimization across entire sequences, inadvertently diluting the impact of critical segments by averaging gradients from both optimal and suboptimal regions.
+This indiscriminate updating forces such methods to rely on larger datasets (4–5× more samples in ablation studies) to statistically filter irrelevant signals and achieve comparable error reduction.
+Furthermore, the utterance-level optimization strategy risks destabilizing well-trained segments through unnecessary perturbations, whereas FPO’s surgical precision preserves existing capabilities while efficiently resolving localized failures.
+TABLE IV: Comparison between our proposed FPO and other preference alignment approaches.
+SDPO UNO RIO Mask-DPO FPO
+Grain Utterance Utterance Utterance Sentence Token
+Goal Overall quality Uncertainty perceptual Robustness enhancement Factuality alignment Targeted correction
+Annotation Global preference Global preference Global preference Global preference Token-level preference
+Reward Binary Binary Binary Binary Fine-grained
+VI Limitations and Discussion
 One of the main limitations of this study lies in the construction of the preference training dataset.
@@
 Although our proposed approach is data efficient and need less annotation data compared with previous works, the fine-grained annotation at each sample is more difficult than utterance level annotation.
-In future work, we plan to explore strategies to alleviate this limitation.
-For example, leveraging weakly supervised, semi-supervised, or self-supervised learning techniques may help reduce the dependence on manually annotated data.
-Ethics Statement
-This work does not raise any ethical concerns.
-All data and pre-trained models used in this study are publicly available and comply with the following licenses: Creative Commons BY 4.0, Creative Commons CC0, Creative Commons BY-NC-ND 4.0, and Creative Commons BY-SA 4.0.
-Appendix A Appendix
-A.1 Segmental Errors
-Temporal modeling errors stem from the model’s limitations in capturing the temporal dynamics of speech, leading to issues such as mispronunciation, where abrupt or unstable pronunciation changes occur due to insufficient modeling of pitch continuity or variation trends.
-Another common problem is abnormal silence, where the model erroneously inserts silent frames or extends silent segments, disrupting the naturalness of generated speech.
-Additionally, the generated speech may exhibit unnatural prosody (i.g.
-unnatural pause), as the model may fail to adequately capture local rhythmic patterns, including stress, duration, and speaking rate, resulting in speech that deviates from natural prosody.
-Semantic-phonetic alignment errors arise from mapping between the phonetic input and the acoustic output.
-These errors often manifest as repetition, where the model over-relies on its previous outputs and redundantly generates the same words or phonemes.
-Alternatively, truncation errors may occur when the language model fails to encode key phonetic information from the input or neglects certain critical elements, leading to synthesized speech that lacks completeness and accuracy.
-A.2 Prosody Annotation
-For prosody issues, we primarily focus on abnormal pauses and annotate fine-grained prosody errors through a three-step process.
-First, we use an automatic speech recognition system to transcribe the speech into text.
-Next, we employ text LLMs to determine ideal pauses in the text and insert break tags.
-Finally, professional annotators evaluate the speech samples based on the break tags, identifying preferred and dispreferred prosody samples.
-Fine-grained timestamps are then obtained using force-alignment tools based on the text and break tags.
+Furthermore, our proposed approach is not inherently restricted to autoregressive TTS models.
+Beyond the evaluated autoregressive backbones, the framework can be naturally extended to discrete non-autoregressive models.
+In these cases, the underlying mechanism of preference-based optimization remains directly applicable without substantial modification.
+For TTS paradigms that rely on continuous representations, the core idea of our method is still conceptually applicable.
+However, applying the method in this setting introduces additional challenges.
+In particular, constructing sufficiently diverse and well-controlled paired samples from continuous representations for effective reinforcement learning becomes non-trivial.
+Meanwhile, practical challenges may arise due to differences in training objectives and generation mechanisms, which we leave for future work.
+VII Conclusion
+In this study, we propose FPO, a novel RL framework that introduces a new paradigm for enhancing zero-shot TTS robustness through localized error correction.
+By systematically categorizing generation errors into temporal modeling discrepancies and semantic-phonetic misalignments, and applying token-level preference optimization, FPO adopts a segment-focused strategy that significantly outperforms baseline systems.
+This validates our hypothesis that targeted intervention on error-prone segments leads to more efficient and effective error correction.
+FPO reduces the bad case ratio by 12% compared to base models through fine-grained perceptual alignment.
+Additionally, it achieves comparable performance using fewer training samples than baseline systems, demonstrating strong data efficiency.
+Experimental results confirm that FPO substantially improves the robustness of the base TTS system, while subjective evaluations further support FPO’s superiority in naturalness and alignment with human perceptual preferences.